OpenStudy (anonymous):

What does the expression i^100+i^101+i^102 equal

OpenStudy (anonymous):

\[\iota ^{100}+\iota ^{101}+\iota ^{102}=\iota ^{100}\left( 1+\iota+\iota ^{2} \right)=\left( \iota ^{2} \right)^{50} \left( 1+\iota-1 \right)\] =?

OpenStudy (anonymous):

are those variables,or are they \(i\) and in \(i=\sqrt{-1}\)?

OpenStudy (anonymous):

the i in i ...

sammixboo (sammixboo):

Do you have choices?

OpenStudy (anonymous):

\[\iota ^{2}=-1\]

OpenStudy (anonymous):

a = i b = -i c = -1 d = 1

sammixboo (sammixboo):

It's A

OpenStudy (anonymous):

\[\left( -1 \right)^{50}\iota =?\]

OpenStudy (anonymous):

thankyou

OpenStudy (anonymous):

yw

sammixboo (sammixboo):

I understand your equation @surhithayer , but I got I which is A, but I am not so sure so he would be more right than me >.<

sammixboo (sammixboo):

@surjithayer

OpenStudy (anonymous):

you are correct.